### v2

# Unified Multimodal Interleaved Document Representation for Retrieval

Abstract
Information Retrieval (IR) methods aim to identify documents relevant to a query, which have been widely applied in various natural language tasks. However, existing approaches typically consider only the textual content within documents, overlooking the fact that documents can contain multiple modalities, including images and tables. Also, they often segment each long document into multiple discrete passages for embedding, which prevents them from capturing the overall document context and interactions between paragraphs. To address these two challenges, we propose a method that holistically embeds documents interleaved with multiple modalities by leveraging the capability of recent vision-language models that enable the processing and integration of text, images, and tables into a unified format and representation. Moreover, to mitigate the information loss from segmenting documents into passages, instead of representing and retrieving passages individually, we further merge the representations of segmented passages into one single document representation, while we additionally introduce a reranking strategy to decouple and identify the relevant passage within the document if necessary. Then, through extensive experiments on diverse IR scenarios considering both the textual and multimodal queries, we show that our approach substantially outperforms relevant baselines, thanks to the consideration of the multimodal information within documents.
Unified Multimodal Interleaved Document Representation for Retrieval
Jaewoo Lee1,2∗   Joonho Ko1∗   Jinheon Baek1∗   Soyeong Jeong1   Sung Ju Hwang1,2 KAIST1 DeepAuto2 jwlee8877@gmail.com, {joonho.ko, jinheon.baek, starsuzi}@kaist.ac.kr
**footnotetext: Equal contribution
1 Introduction
Information Retrieval (IR) is the task of fetching relevant documents from a large corpus in response to an input query, which plays a critical role in various real-world applications including web search engines and question-answering systems (Shah et al., 2019; Lewis et al., 2020; Guu et al., 2020). Over the years, IR methods have evolved significantly, with approaches broadly categorized into sparse and dense retrieval paradigms. Specifically, sparse retrieval methods Robertson et al. (1994); Jones (2004) focus on lexical overlap between queries and documents; meanwhile, dense retrieval methods Karpukhin et al. (2020); Xiong et al. (2021) utilize neural embeddings to represent queries and documents in a continuous vector space. Note that, recently, dense retrieval methods have gained more popularity over sparse methods due to their capability to capture semantic nuances and context beyond simple keyword matching, leading to multiple successes with improved performance.
Figure 1: Comparison of different IR approaches. (a): Conventional methods use a small portion of the text within the document for its representation. (b): Recent methods use first-page screenshot images to represent the document. (c): Our approach leverages the full contextual information within documents interleaved with multiple modalities by considering them in their original format, and is further capable of pinpointing relevant sections for the query.
Despite their huge successes, existing (dense) retrieval methods face a couple of severe challenges. First, they primarily rely on the textual data for document embedding and retrieval, overlooking the fact that modern documents often contain multimodal content, such as images and tables (beyond the plain text), which can carry critical information that may be essential for accurately understanding and retrieving the relevant documents. To be specific, a diagram within a medical article can more effectively represent the structure of a molecule or the progression of a disease, offering more clarity that would be difficult to achieve with text alone, and omitting such multimodal content can lead to an incomplete understanding (and potentially inaccurate retrieval) of the documents. Also, the segmentation of long documents into discrete passages, which is commonly employed by existing retrieval models to handle the length limitation for embeddings, may prevent models from capturing the full context and the intricate relationships between different parts of the document, ultimately leading to suboptimal retrieval performance. It is worthwhile noting that, concurrent to our work, while there has been recent work that screen captures the document and then embed its screenshots (to consider different modalities in a unified format) (Faysse et al., 2024; Ma et al., 2024), not only its content (such as paragraphs, images, and tables) can be fragmented into different sub-images, leading to the loss of contextual coherence across the entire document, but also the visual representation of text may hinder the model’s ability to capture the semantic relationships present in the original textual data, while increasing the image resolution leads to the concern on the memory requirements.
In this work, we introduce a novel approach to holistically represent documents for IR, which addresses the aforementioned challenges by representing and retrieving the documents interleaved with multiple modalities in a unified manner (See Figure 1). Specifically, our method revolves around the recent advance of Vision-Language Models (VLMs), which enable the processing and integration of multimodal content (such as text, images, and tables) directly into a single token sequence, thereby preserving the context and relationships between various parts of the document, unlike the previous approaches that rely on the fragmented visual representations. Also, in cases where the number of tokens in a document is large and exceeds the capacity of a single context window of VLMs, we propose a strategy to segment the document into passages, each represented within the token limit, and combine these passage embeddings into a unified document representation. This strategy differs from existing IR approaches that independently represent and retrieve at the passage level, potentially losing the overall document context. Lastly, to accurately identify only the relevant sections within the retrieved lengthy document, we introduce a reranking mechanism that is trained to pinpoint the passage most pertinent to the query (among all the other passages within the document), allowing for both the coarse-grained document-level matching and fine-grained passage-level retrieval. We refer to our overall framework as Interleaved Document Information Retrieval System (IDentIfy).
We experimentally validate the effectiveness of IDentIfy on four benchmark datasets, considering both the text-only and multimodal queries. On a battery of tests conducted, we observe that our approach substantially outperforms relevant baselines that consider only the uni-modality for document representations, thanks to the consideration of multimodal content. Further, we find that the strategy to represent the whole document with its single representation (by merging embeddings of its splits) is superior to the approach of individually representing them for document retrieval, but also performing reranking over the sections of the retrieved document is superior to the approach of directly retrieving those sections, which confirm the efficacy of the proposed retrieval and reranking pipeline for document and passage retrieval, respectively.
2 Related Work
Information Retrieval
Information Retrieval (IR) involves finding documents relevant to a query, which plays a crucial role in applications such as search and question-answering (Zhu et al., 2023; Gao et al., 2023; Ram et al., 2023; Shi et al., 2024; Jeong et al., 2024a). Earlier IR approaches measured the similarity between queries and documents based on their lexical term matching, such as BM25 and TF-IDF (Robertson et al., 1994; Jones, 2004). Yet, these methods often struggled to capture the semantic nuances beyond surface-level term overlaps. To overcome this, along with advancements in language models (Devlin et al., 2019; Liu et al., 2019), there have been dense retrieval approaches that embed both the queries and documents into a shared dense vector space (Karpukhin et al., 2020; Xiong et al., 2021), enabling the calculation of semantic similarity between them more effectively by capturing the deeper contextual information. Yet, previous studies have mainly focused on enhancing the textual representations of queries and documents, while overlooking the multimodal nature of documents beyond text, which can potentially provide richer context and aid in more accurate retrieval (Liu et al., 2021; Jeong et al., 2024b).
Multimodal Information Retrieval
Recent studies in IR have expanded the focus from purely text-based retrieval models to those that consider other modalities, such as images (Radford et al., 2021; Xiao et al., 2024), tables (Herzig et al., 2021; Chen et al., 2024) and graphs (Baek et al., 2023); however, the majority of these approaches (Zhou et al., 2024; Long et al., 2024; Lerner et al., 2024; Nowak et al., 2024; Caffagni et al., 2024) have primarily explored how to process the multimodal queries, meanwhile, they often overlook the equally important multimodal characteristics of the documents being retrieved. In efforts to handle diverse multimodal elements within documents, there are concurrent studies that have proposed to capture screenshots of documents, such as PDFs (Faysse et al., 2024; Cho et al., 2024) or Wikipedia web pages (Ma et al., 2024), and subsequently encoding them through vision models (Ding et al., 2024). However, these methods are not only limited by factors, such as image resolution and computational memory, constraining their application to documents longer than a single page [^1] For instance, Ma et al. (2024) requires processing 9.8k image tokens just to process a single-page document, and it results in 2TB of storage for handling the entire Wikipedia corpus, which may not be practical., but also fall short by treating the diverse modalities within a document as a single visual entity, leading to suboptimal document representations that may fail to effectively capture the nuanced interdependence between text and images. Also, while there are concurrent studies Jiang et al. (2024b); Lin et al. (2024) that consider images and text as retrieval targets, they primarily focus on representing image-text pairs and their retrieval, rather than addressing the holistic representation of documents that include multiple images and another modality (tables). Finally, all the aforementioned work does not address the issue of splitting documents into smaller fragments (passages or sub-images), which may disrupt the holistic contextual view of the entire document.
Vision-Language Models
Recently developed Vision-Language Models (VLMs) have emerged as a powerful tool for jointly processing visual and textual data, which combine the image understanding capabilities of visual encoders (Radford et al., 2021; Zhai et al., 2023) with the advanced reasoning abilities of language models (OpenAI, 2022, 2023a). These models have achieved remarkable performance across diverse vision-language (VL) tasks (such as image captioning and visual question answering) (Dai et al., 2023; OpenAI, 2023b), with the substantially limited attention on their applications to IR. We note that the latest developments in this field have particularly focused on enabling VLMs to handle interleaved, multimodal content, which involves a mixed sequence of images and text (Zhang et al., 2023; Li et al., 2024b). In particular, LLaVA-NeXT-Interleave (Li et al., 2024b) introduces a fine-tuning approach that specifically enhances the VLMs’ capacity to understand complex interleavings of multiple images and text within a single context. Drawing inspiration from these advances, we propose to harness the capabilities of VLMs to create unified embeddings for documents interleaved with text and images (as well as tables).
3 Method
We present IDentIfy to holistically represent documents interleaved with multimodal elements.
3.1 Preliminaries
We begin with preliminaries, formally explaining information retrieval and vision-language models.
Information Retrieval
IR is formally defined as the task of identifying a set of relevant documents $\{{\bm{d}}_{1},{\bm{d}}_{2},\ldots,{\bm{d}}_{k}\}\subseteq\mathcal{D}$ from a large corpus $\mathcal{D}$, given an input query ${\bm{q}}$. Here, each query ${\bm{q}}$ and document ${\bm{d}}$ are represented as a sequence of tokens: ${\bm{q}}=[q_{1},q_{2},\ldots,q_{n}]$ and ${\bm{d}}=[d_{1},d_{2},\ldots,d_{m}]$, and traditional IR approaches typically consider these tokens as purely textual elements. However, we propose to extend this assumption to have the tokens of both the textual and visual content, to capture the multimodal nature of many real-world documents. Then, this new extension raises important questions of how can both the textual and visual content be represented within a unified token framework, and how can these multimodal tokens be seamlessly integrated and encoded for document representations. To answer them, we harness the power of recent vision-language models below.
Figure 2: Overview of the proposed IDentIfy. (a): In our document retriever, a query encoder represents a query (purple), and sections are encoded with a section encoder whose embeddings averaged to form a document representation (blue). Contrastive learning loss (red) is used for training the document retriever. (b): Reranker scores query-section relevance with the concatenation of the query and section, trained using Binary Cross-Entropy loss.
Vision-Language Models
We now turn to describing Vision-Language Models (VLMs), which are designed to jointly encode the textual and visual information in a unified token framework. These models are generally comprised of two main components: a visual encoder and a language model, interconnected through a projection layer. Specifically, given the document that may contain interleaved modalities (e.g., text and images), the visual encoder extracts high-level visual features from (multiple) images embedded within the document, mapping them into a latent space. Then, these visual features are transformed into a sequence of visual tokens via the projection layer, represented as follows: $\mathbf{V}\!\in\!\mathbb{R}^{V\!\times\!d_{\text{emb}}}$, where $V$ denotes the visual token length and $d_{\text{emb}}$ is the token dimension size. Similarly, for the textual content embedded within the document, the language model uses a word embedding layer to convert the input text into a sequence of tokens, as follows: $\mathbf{L}\!\in\!\mathbb{R}^{L\!\times\!d_{\text{emb}}}$, where $L$ denotes the token length of text.
In this work, we also propose to account for tables that are an integral modality for holistically representing the full content of documents. However, in contrast to text and images that have dedicated processing layers within the VLM architectures, tables do not have a specific representation layer. Nevertheless, we argue that recent VLMs are pre-trained on diverse web data, and subsequently they are implicitly learned to handle the table structures formatted in HTML. Consequently, we treat HTML-format table data as a linearized sequence of HTML words, applying the same word embedding layer as is used for plain text. To be formal, this process converts the table content into table tokens, as follows: $\mathbf{T}\!\in\mathbb{R}^{T\!\times\!d_{\text{emb}}}$, where $T$ is the token length of the table. Lastly, once extracted, the visual tokens, text tokens, and table tokens are concatenated (to form a unified token sequence) and then passed through the remaining layers of VLMs, to capture both uni- and cross-modal relationships across different modalities, ultimately enabling the comprehensive understanding of the documents.
3.2 Retriever
We now explain how we design a retriever specifically tailored for multimodal interleaved document retrieval. In particular, our approach leverages a VLM capable of processing text, images, and tables within a single document. Further, following the standard practice of existing retrieval architectures (Karpukhin et al., 2020; Xiong et al., 2021), we use a dual-encoder structure, which consists of a query encoder and a section encoder, both are based on the VLM, illustrated in Figure 2 (a).
[…]
The remaining step to discuss here is how to train those two query and section encoders for IR. Recall that the goal of the retriever is to assess a relevance score between the query and the document. To achieve this goal, we use a contrastive learning loss based upon the query and document representations, whose objective is to assign higher similarity scores to relevant documents (positive samples) and lower scores to irrelevant ones (negative samples) for the query, formulated as follows:
$\mathcal{L}_{\text{retriever}}\!$ $=\!-\frac{1}{B}\sum^{B}_{i=1}\log\left(\frac{\phi\left(\mathbf{Z}_{\mathrm{Q}_{i}},\mathbf{Z}_{\mathrm{D}_{i}}\right)}{\sum^{B}_{j=1}\phi\left(\mathbf{Z}_{\mathrm{Q}_{i}},\mathbf{Z}_{\mathrm{D}_{j}}\right)}\right),$
$\phi\left(\mathbf{a},\;\mathbf{b}\right)\!$ $=\!\exp\left(\frac{\mathbf{a}^{\top}\mathbf{b}}{\lVert\mathbf{a}\rVert\lVert\mathbf{b}\rVert}\right),$ (1)
where $B$ is the batch size during the training phase. Here, by minimizing $\mathcal{L}_{\text{retriever}}$, the retriever learns to optimize the similarity between queries and their relevant documents, enabling the retrieval of the most pertinent documents (among all) for the given input query during inference.
3.3 Reranker
To enable fine-grained retrieval within documents beyond the retrieval of documents themselves, we introduce a section-level reranking mechanism that identifies the section most relevant to the query. In particular, once the document is retrieved, the objective of the reranker $f_{\mathrm{R}}$ is to pinpoint the specific sections within the document that best match the query. We also note that this reranker is similarly operationalized with the VLM along with a binary classifier on top of it, which directly measures the relevance of each query-section pair (Figure 2 (b)).
[…]
To train this reranker, we use the binary cross-entropy loss, formulated as follow:
$\mathcal{L}_{\text{reranker}}$ $\!=\!\sum_{i=1}^{B}\sum_{j=1}^{S_{i}}\!\frac{1}{BS_{i}}\ell\left(\mathbf{y}_{{\bm{s}}_{i,j}},\;f_{\mathrm{R}}\left(\left[{\bm{q}},\;\hat{{\bm{s}}}_{i,j}\right]\right)\right),$
$\ell\left(y,\hat{y}\right)$ $\!=\!-\left[y\log\hat{y}\!+\!(1\!-\!y)\log(1\!-\!\hat{y})\right],$ (2)
[…]
4 Experiments
Table 1: Results with different document formats for retrieval.
Format R@1 R@10 R@100 MRR@10
Entity 3.1 15.5 39.7 6.1
Summary 13.4 41.3 66.5 21.6
Text-document 12.5 37.8 68.7 19.8
+ Single-image 16.4 45.4 77.1 25.3
+ Interleaved (Ours) 20.5 50.0 78.0 29.4
Table 2: Results with different section retrieval strategies. Document (Ours) performs document retrieval and section reranking, whereas Passage performs section retrieval and reranking. * denotes the model without reranking.
Granularity R@1 R@10 R@20 MRR@10
Passage* 3.9 16.9 22.0 7.5
Passage 28.6 36.4 37.8 31.2
Document (Ours) 35.1 50.8 53.6 40.3
Table 3: Performance on document retrievals. (a): Results of document retrieval for multimodal queries on InfoSeek and ViQuAE. (b): Results of document retrieval for textual queries on Encyclopedic-VQA (Enc-VQA) and ViQuAE.
(a) Document Retrieval with Multimodal Queries
Format Dataset R@1 R@10 R@100 MRR@10 Text-document InfoSeek 6.8 23.6 52.5 11.2 + Interleaved 10.2 30.4 57.3 15.7 Text-document ViQuAE 13.5 40.4 67.4 20.9 + Interleaved 17.5 46.0 69.4 26.3
(b) Document Retrieval with Textual Queries
Format Dataset R@1 R@10 R@100 MRR@10 Text-document Enc-VQA 62.7 76.3 87.4 67.0 + Interleaved 65.4 76.8 87.8 69.0 Text-document ViQuAE 55.8 71.5 83.0 60.9 + Interleaved 56.5 72.2 83.0 61.6
4.1 Experimental Setups
Datasets
We evaluate IDentIfy on four benchmark datasets designed for multimodal IR tasks that require understanding of both the textual and visual cues within queries and documents, as follows: Encyclopedic-VQA (Mensink et al., 2023) is a large-scale benchmark for multimodal Visual Question Answering (VQA) with queries linked to specific Wikipedia sections and includes both text-only and multimodal queries; InfoSeek (Chen et al., 2023) is a knowledge-intensive VQA dataset with multimodal questions generated from Wikidata triples that include diverse entities such as landmarks, animals, and food; ViQuAE (Lerner et al., 2022) involves both text-based and multimodal queries about human entities, linked to annotated Wikipedia sections, making it ideal for evaluating section retrieval; Open-WikiTable (Kweon et al., 2023) extends WikiSQL (Zhong et al., 2017) and WikiTableQuestions (Pasupat and Liang, 2015), targeting open-domain table QA by identifying documents or sections containing relevant tables. We provide more details on datasets in Appendix A.
Baselines
We compare our approach against diverse baselines that capture different document representations. First, the Entity and Summary baselines retrieve documents based on their titles and summary sections, respectively, leveraging high-level textual cues. Also, the Text-document retriever baseline utilizes the full textual content of documents for retrieval. We further include the Single-image baseline that additionally leverages the first image of each document. IDentIfy is our model that holistically represents multimodal content (text, images, and tables) within documents.
Evaluation Metrics
To evaluate our approach, we use standard metrics: Recall@K (R@K) measures whether the relevant document or section appears within the top-K results; MRR@K measures how early the first relevant item is ranked (within top-K) by averaging its inverse rank across queries.
Implementation Details
We use LLaVA-NeXT-Interleave (Li et al., 2024b) of 0.5B parameters as the basis VLM for both the retriever and reranker. During training, documents are represented using randomly selected four sections, while in inference, we consider all sections within each document. For section-level retrieval, all sections within the top 25 retrieved documents are reranked. Experiments are conducted on a single H100 GPU.
4.2 Experimental Results and Analyses
Main Results
We report retrieval performance on the Encyclopedic-VQA dataset in Table 2, where queries include both text and images. We observe that IDentIfy achieves the best performance, improving R@1 scores by 53.0%, 64.0%, and 25.0% over Summary, Text-document, and Single-image retrieval baselines, respectively, with similar trends observed for other metrics. These results demonstrate the effectiveness of integrating multimodal content holistically into a unified representation. To further illustrate the advantages of our approach, we provide case studies in Appendix C.
We further examine the impact of our document retrieval and section reranking pipeline. In Table 2, the passage retriever represents individual sections as separate retrieval units, whereas the document retriever (ours) aggregates multiple section representations into a single representation. Then, we perform reranking over the retrieved sections or the sections from the retrieved documents, and then report the results in Table 2 (where * denotes the model without reranking). From this, we observe that the passage retriever without reranking (Passage*) achieves suboptimal retrieval performance, highlighting the challenge in pinpointing the most relevant section within a document using traditional retrieval methods. In contrast, when the reranker is used alongside the document retriever, the performance significantly surpasses that of the passage retrieval. These results confirm the importance of leveraging holistic context from multiple, interrelated sections within documents.
Interleaved format enhances document retrieval across modalities.
Table 4: Performance on section reranking. (a): Results of section reranking for multimodal queries on Encyclopedic-VQA (Enc-VQA) and ViQuAE. (b): Results of section reranking for textual queries on Enc-VQA and ViQuAE. For fair comparison between different methods, the reranking target (i.e., candidate sections from retrieved documents) are the same across them.
(a) Section Reranking with Multimodal Queries
Format Dataset R@1 R@10 R@20 MRR@10 Text-document Enc-VQA 40.7 52.8 55.5 44.8 + Interleaved 42.4 53.6 55.7 46.3 Text-document ViQuAE 12.6 31.7 37.7 18.2 + Interleaved 11.4 32.1 39.2 17.5
(b) Section Reranking with Textual Queries
Format Dataset R@1 R@10 R@20 MRR@10 Text-document Enc-VQA 68.1 79.4 80.2 72.3 + Interleaved 69.7 80.1 80.6 73.6 Text-document ViQuAE 27.8 50.2 57.7 35.0 + Interleaved 29.9 50.9 59.8 36.7
Table 5: Retrieval results for tables, where Zero-shot denotes a model trained on Encyclopedic-VQA but not on the target dataset. Finetuned refers to additional training of the model on the target dataset. (a): Results for tabular document retrieval on Open-WikiTable (OWT). (b): Textual and tablular section retrieval results on ViQuAE and OWT datasets, respectively. (c): Reranker accuracy of a classification task that identifies the section containing the query-associated table given a gold document.
(a) Document Retrieval for Tables
Model R@1 R@10 R@100 MRR@10 Zero-shot 29.4 58.0 86.0 38.1 Finetuned 55.8 84.1 93.5 66.1
(c) Tabular Classification
Model Random Zero-shot Finetuned Acc@1 11.9 9.3 56.5
(b) Section Retrieval for Tables
Model Modality Dataset R@1 R@10 R@20 MRR@10 Zero-shot Text ViQuAE 20.3 49.0 57.7 28.9 Finetuned 29.9 50.9 59.8 36.7 Zero-shot Table OWT 5.9 20.5 29.4 9.1 Finetuned 8.4 36.7 52.8 15.2
We further expand our experiments to two additional datasets, InfoSeek and ViQuAE, and report document retrieval results. As shown in Table 3, our model consistently outperforms the Text-document baseline for both multimodal and text-only queries. We attribute these gains to the integration of multimodal content, allowing the VLM to capture richer alignments and leverage pre-existing knowledge for more effective document representation (Xu et al., 2024).
Interleaved format is also beneficial in section retrieval.
Similarly, we evaluate section retrieval performance on Encyclopedic-VQA and ViQuAE datasets, for both multimodal and textual queries. As shown in Table 4, our model outperforms the Text-document baseline in most cases. However, the performance gains over the baseline are smaller compared to the document retrieval setup. This is likely because section reranking focuses on evaluating the relationship between a single section and a query (rather than leveraging the holistic context of the entire document), and individual sections may lack the diverse multimodal information necessary for fully capturing the intent of queries.
Retrieving tables interleaved within documents is challenging.
We explore the retrieval task for tabular data, aiming to identify documents or sections containing query-relevant tables, and compare models trained on Encyclopedic-VQA (Zero-shot) with those additionally trained on Open-WikiTable (Finetuned). As shown in Table 5 (a), the Finetuned retriever outperforms the Zero-shot retriever on retrieving documents containing query-relevant tables. However, more fine-grained section reranking results (identifying sections containing query-relevant tables) in Table 5 (b) may reveal a notable modality-specific challenge: the performance of Zero-shot and Finetuned rerankers is considerably lower on table retrieval compared to their performance on text retrieval, despite both the text and tables being represented with word tokens. To better understand this, we design a classification task, where rerankers are tasked with identifying the correct section containing the target table within the golden document. Then, as shown in Table 5 (c), the Zero-shot reranker performs comparably to random selection, while the Finetuned reranker shows modest improvements. These findings highlight the intrinsic challenge of tabular retrieval, suggesting the need for table-specific modules to more holistically represent multimodal interleaved documents.
More sections enhance document retrieval performance but raise computational costs.
Figure 3: Trade-off between performance (MRR@10) and training cost (GPU Memory) for retrieval.
Table 6: Comparison of negative sample selection strategies for reranker training: Top-K (top-k retrieved sections), In-batch (sections from other samples in the batch), and In-document (sections in the same document).
Negative R@1 R@20 MRR@10
Top-K 38.1 55.3 44.4
In-batch 39.5 55.4 45.0
In-document (Ours) 42.4 55.7 46.3
Table 7: Comparison of different training objectives for the reranker: Contrastive considers sections as retrieval units and uses the one for document retriever training; Document + BCE concatenates the query with multiple sections from the same document and uses the BCL loss; Section + BCE trains the reranker by concatenating the query with each section individually.
(a) Section Retrieval for Multimodal Queries
Train Loss R@1 R@10 R@20 MRR@10 Contrastive 3.6 15.0 21.3 6.5 Document + BCE 13.6 29.6 32.9 24.1 Section + BCE (Ours) 42.4 53.6 55.7 46.3
(b) Section Retrieval for Textual Queries
Train Loss R@1 R@10 R@20 MRR@10 Contrastive 13.6 37.7 45.1 20.6 Document + BCE 23.8 43.4 47.2 39.1 Section + BCE (Ours) 69.7 80.1 80.6 73.6
To see how the number of sections used for representing each document impacts performance, we evaluate document retrieval on the InfoSeek dataset by varying the sections per document during training. As shown in Section 4.2, incorporating more sections improves MRR@10 from 7.5 to 15.7 due to leveraging richer multimodal and contextual information. However, this comes at the cost of increased computational requirements, as processing more sections raises GPU memory consumption.
Sections from the same document act as effective negatives to enhance reranker performance.
In training the reranker, we investigate whether considering sections from the same document as negative examples (called In-document) is effective than other strategies, such as Top-K negatives (top-K retrieved sections based on their similarity with the input query) and In-batch negatives (positive sections from other samples in the same batch). As shown in Section 4.2, we observe that the In-document approach achieves superior performance especially on R@1, demonstrating its ability to effectively identify the most pertinent section among highly similar sections within the same document, i.e., its training objective can encourage the reranker to focus on fine-grained distinctions between closely related sections (within the same document).
BCE loss is the most effective to train the section reranker.
In our reranker design, we use a binary cross-entropy (BCE) loss by concatenating the query with each document section individually (Section + BCE), allowing the model to directly assess query-section relevance. As an alternative, we also explore a contrastive loss (Contrastive), which models section reranking similarly to document retrieval but uses sections as the retrieval units, and a variant of BCE loss (Document + BCE), where the query is concatenated with multiple sections (both positive and negative) from the same document. As shown in Table 7, the Section + BCE reranker outperforms both alternatives. Specifically, contrastive loss performs the worst, suggesting that direct concatenation of query and section provides clearer relevance signals, consistent with conventional reranking approaches. Moreover, while Document + BCE leverages inter-section context, its performance might be hindered by training constraints as the model processes fewer sections during training (Jiang et al., 2024a; Lee et al., 2024), and addressing it would be interesting future work.
Our Approach is Robust with Different Base Models.
Table 8: Results with another base model (LLaVA-OneVision) for document retrieval (with different document formats).
Format R@1 R@10 R@100 MRR@10
Entity 2.3 10.3 29.7 4.3
Summary 7.6 24.7 55.7 12.0
Text-document 7.0 24.1 50.4 11.7
+ Single-image 9.3 31.4 61.9 15.4
+ Interleaved (Ours) 12.1 36.1 62.5 18.2
To ensure the effectiveness of our approach across VLMs, we evalulate its performance with another VLM, LLaVA-OneVision (Li et al., 2024a), with 0.5 billion paramters, in addition to LLaVA-NeXT-Interleave (Li et al., 2024b) used in our main experiments. Results in Table 8 show that ours continues to outperform baselines, achieving a notable 30.1% gain in R@1 over the best baseline.
5 Conclusion
In this paper, we introduced IDentIfy, a novel IR framework designed to address the limitations of conventional methods that rely on textual content of documents and their segmented passages. Specifically, our approach sits on top of recent VLMs, which enables integration and representation of diverse multimodal content (including text, images, and tables) into a unified document representation. Also, unlike previous strategies that segment documents at the passage level, our method merges these segments to maintain the document’s structural coherence, while further introducing a reranking strategy for precise identification of relevant sections. Extensive experiments across various IR datasets demonstrated that IDentIfy consistently outperforms existing baselines, confirming that the interleaved multimodal representation significantly enhances the quality of the document retrieval. We believe IDentIfy represents a crucial step toward more comprehensive and contextually aware IR systems, capable of handling the increasing multimodality of modern information sources.
Limitations
Due to the limitations of a single H100 GPU, we represent documents by selecting a limited number of sections and averaging their corresponding embeddings. While this reduces the computational demands, our findings suggest that capturing a broader document context leads to improved retrieval performance. Hence, leveraging the long context window of LVLMs could further enhance document retrieval by capturing more comprehensive information from the full document. Moreover, our reranker design follows the conventional approach of concatenating the input query with individual sections. However, we believe that providing the reranker with all the sections together would allow the model to better leverage the contextual information from the entire interleaved document, potentially resulting in improved performance. In order to fully leverage the interleaved format in the IR system, addressing the issues by reducing the GPU load when processing interleaved documents would greatly boost overall IR performance. We leave these explorations for future work.
Ethics Statement
In this work, we use a publicly available retrieval corpus for information retrieval tasks. However, the retrieval corpus may contain private, harmful, or biased content. Such undesirable features could unintentionally be reflected in the behavior of retrievers and rerankers trained on this data, potentially leading to ethical concerns during real-world deployment. However, current information retrieval techniques, including ours, do not address the retrieval of undesirable content. We recognize the critical need for safeguards to mitigate this issue. This is essential to ensure that information retrieval systems are reliable, fair, and safe for deployment.
Appendix A Details of Experimental Setups
Dataset configuration
Table 9 summarizes the key properties of the datasets used in our experiment, including query modality, target item, entity domain, number of entities, and whether a section ID is provided to indicate the section containing the answer. Additionally, we provide the number of samples in the training, evaluation, and test splits, as well as the size of the corpus. We provide a more detailed explanation of the datasets below.
  • Encyclopedic-VQA (Mensink et al., 2023) is a large-scale visual question-answering (VQA) benchmark dataset, widely used for measuring the performance of multimodal IR models. Each query is linked to a specific section of a Wikipedia document (containing an answer for it) and is manually annotated by humans. Also, this dataset offers both text-only and multimodal queries. In addition to this, the queries are related to fine-grained properties of species and landmarks. Our experiments focus on the single-hop category where questions can be answered in a single retrieval step.
  • InfoSeek (Chen et al., 2023) is a dataset designed for knowledge-intensive VQA, covering a wide range of entities (such as landmarks, animals, and food). Questions are generated by filling human-written templates with knowledge triples (subject, relation, object) available from Wikidata, which involve only the multimodal queries. As the test dataset is not available, we use the validation set as our test set, and split the training set into training and validation subsets with a 9:1 ratio.
  • ViQuAE (Lerner et al., 2022) is a dataset focused about human entities. It provides both textual and multimodal queries, with each query linked to a specific section of a Wikipedia document that contains an answer annotated by humans, which makes it an ideal benchmark for section retrieval.
  • Open-WikiTable (Kweon et al., 2023) is an extension of WikiSQL (Zhong et al., 2017) and WikiTableQuestions (Pasupat and Liang, 2015), designed for open-domain table question answering that requires retrieval of the most relevant table from a broader corpus. For our experiments, we adapt the WikiTableQuestions subset of Open-WikiTable, aiming at identifying the document or document section containing the target table.
Dataset pre-processing
In our study, we leverage interleaved multimodal content from Wikipedia documents. However, existing corpora associated with IR datasets often lack this content, typically only including the first few words of each document. Therefore, we download the HTML file of each Wikipedia document for corpus augmentation.
If the dataset provides Wikipedia URLs for its corpus, we use them to download the HTML files. Alternatively, if only entity names are provided, we generate Wikipedia URLs using those names. If a Wikipedia URL is deprecated, we remove the corresponding document from the corpus along with any associated queries. From the HTML files, we extract text, image URLs, and tables. We then split the contents by subtitles in the document where each chunk corresponds to a section. For the images, we use the image URLs to download the corresponding images, removing any invalid URLs. This process produces a dictionary that organizes text, images, and tables by section.
Since downloading contents for all documents across datasets is time- and memory-intensive, we preprocess subsets of each corpus, including documents relevant to queries in the training, evaluation, and test splits, along with unrelated documents.
Implementation Details
To take advantage of larger batch sizes (while reducing GPU memory usage), we apply LoRA (Hu et al., 2022). Also, to further optimize the GPU usage, we scale each image down to half of its original height and width and then combine four scaled-down images into a single composite image. All experiments are conducted using a single H100 GPU.
Appendix B Additional Experimental Results
Data Requirements for Models
Refer to captionRefer to caption
(a) Retriever performance
Refer to captionRefer to caption
(b) Reranker performance
Figure 4: Retrieval performance with different dataset sizes for training. (a): When training a retriever, large datasets rather deteriorate the retrieval performance as it may be overfitted, resulting in low generalization. (b): On the other hand, a larger dataset size is beneficial to training a re-ranker.
We analyze the effect of different dataset sizes for training on retriever and reranker performance. To achieve this, we randomly prune samples in the Encyclopedic-VQA dataset at various ratios and report the performance of models trained on these subsets. In Figure 4 (a), we observe that too many samples can degrade retrieval performance. Also, retrieval of textual queries requires fewer samples to reach its optimal performance compared to multimodal retrieval. Similarly, in Figure 4 (b), section retrieval for multimodal queries requires 10% of the dataset to achieve 80% of the full-dataset performance, while section retrieval for textual queries needs only 5%. These observations suggest that additional modalities increase the need for more data. This accounts for the inferior performance of the interleaved format in the ViQuAE experiments (Table 4 (a)). The ViQuAE dataset, at only 2.2% of the size of Encyclopedic-VQA, may be small for the reranker to effectively learn multimodal query-section alignments. We also observe that section retrieval is more challenging, with more samples improving the reranker’s performance. This explains why the ViQuAE reranker has much lower section retrieval scores compared to the one trained on the Encyclopedic-VQA (Table 4 (b)). Given the challenge of obtaining large query-section pair samples, exploring more effective reranker training pipelines is necessary.
Appendix C Case Studies of Document Retrieval
We conduct case studies to demonstrate the advantages of our approach in document retrieval with textual and multimodal queries. In Figure 5 and Figure 6, we illustrate the instances where our approach, which leverages interleaved multimodal contents (e.g., images, tables, and text) within documents, retrieved correct documents for given queries, while the conventional one, which represents documents using only textual data, retrieved documents that appeared to be relevant but were not actually related to the queries.
In Figure 5, a textual query asks for the name of the park located on the north shore of Foster Reservoir. The conventional approach retrieved a document containing unrelated information about a different reservoir. While this document includes terms such as "Peak District National Park" and "North America farm," which make the document superficially relevant, it fails to answer the query. In contrast, our approach identified the document containing the correct answer to the given query.
The advantages of integrating multimodal content into document representation become more apparent in document retrieval with multimodal queries, as shown in Figure 6. For a query consisting of an image of a town hall in Hanover and a textual question about its designer, both our approach and the conventional one retrieved documents about town halls in Germany. However, our approach pinpointed the exact document about the town hall in Hanover, indicating that Hermann Eggert designed the building. The conventional method retrieved a document about a town hall in Munich, which is somewhat related but not an exact match to the query image or question.
These cases underscore the benefits of leveraging multimodal content in information retrieval. Integrating interleaved multimodal elements, our approach aligns more effectively with the input query, resulting in more accurate and fine-grained retrieval. This superiority is supported by Xu et al. (2024), which highlights that models perform better when prompted with rich multimodal information, enabling them to capture alignments across modalities and enhance the representation of given inputs.
Q: What is the name of the park on the north shore of foster reservoir?
(a) Interleaved Multimodal Document Retrieval ,,,,,,,,,,,,, (b) Text-only Document Retrieval
Figure 5: Retrieved documents across different document formats for document retrieval with a given textual query. (a): A document retrieved when represented leveraging interleaved multimodal contents within documents (ours). (b): A document retrieved when using only textual format
[Uncaptioned image] Q: Who designed this building?
(a) Interleaved Multimodal Document Retrieval ,,,,,,,,,,,,, (b) Text-only Document Retrieval
Figure 6: Retrieved documents across different document formats for document retrieval with a given multimodal query. (a): A document retrieved when represented leveraging interleaved multimodal contents within documents (ours). (b): A document retrieved when using only textual format